I have a formula and a problem. please help step by step Thanks
An object that is falling or vertically projected into the air has its height, in feet, above the ground given by s=-16t2+vot=so formula
where s is the height in feet, vo is the original velocity of the object in feet per second, t is the time the object is in motion, in seconds, and so is the height, in feet, from which the object is dropped or projected. A ball is thrown straight up from a rooftop at an original velocity of 80 feet per second from a height of 96 feet. The ball misses the rooftop on its way down and eventually stikes the ground. How high will the ball be 4 seconds after being thrown. The book states 160 feet but how do I put this in formula. Thanks

Answer by ankor@dixie-net.com(22740)

An object that is falling or vertically projected into the air has its height, in feet, above the ground given by s=-16t2+vot=so formula
where s is the height in feet, vo is the original velocity of the object in feet per second, t is the time the object is in motion, in seconds, and so is the height, in feet, from which the object is dropped or projected. A ball is thrown straight up from a rooftop at an original velocity of 80 feet per second from a height of 96 feet. The ball misses the rooftop on its way down and eventually stikes the ground. How high will the ball be 4 seconds after being thrown.
:
The equation with the given initial velocity and height
:
s = -16t^2 + 80t + 96
:
Given: t = 4, substitute 4 for t:
s = -16(4^2) + 80(4) + 96
s = -16(16) + 320 + 96
s = -256 + 320 + 96
s = 160 ft